OpenStudy (anonymous):

Solve M = 2x + 3y for y.

OpenStudy (anonymous):

incomplete question

OpenStudy (anonymous):

nope its not

OpenStudy (lgbasallote):

if you're familiar with MDAS, in these cases you go with SADM so start with subtraction/addition...take away 2x..do this by subtracting both sides by 2x M - 2x = 3y now divide both sides by 3 to isolate y M - 2x = y ----- 3

OpenStudy (anonymous):

ty!!!

OpenStudy (anonymous):

wait M-2 and 2x -2?

OpenStudy (anonymous):

then divide by 3?

OpenStudy (anonymous):

or are u saying this>>>>>>>>> M - 2x = 3y -2 -2

OpenStudy (anonymous):

then divide 3

OpenStudy (anonymous):

?

OpenStudy (lgbasallote):

M = 2x + 3y -2x -2x M - 2x = 3y then divide by 3

OpenStudy (anonymous):

k ty!!

OpenStudy (lgbasallote):

welcome